OpenStudy (anonymous):

If energy is conserved, then: A.PE(before) = KE(before). B.initial (PE + KE) = final (PE + KE). C.the momentum does not change. D.the initial KE must be zero.

OpenStudy (anonymous):

if you have a book up in the shelf, it has PE1=mgh and KE1=0.5*m*v^2=0 (doesn't have velocity). If it drops, on the moment it hits the floor PE2=mgh=0 and KE2=0.5*m*v^2. Where PE1+KE1=PE2+KE2.

OpenStudy (anonymous):

b.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes you are right :)